What is the market value of a bond that has a 6 % coupon paid semiannually , has a $ 2,000 face value , has 15 years to maturity and has a 4 % yield to maturity . Write down the formula with appropriate numbers inserted to solve the problem along with the correct answer .
Answer #1

Face value of Bond (F) = $2,000

Coupon rate = 6%

Coupon frequency = semiannually

Semiannual Coupon Amount(C) = 3%*2000= $60

Semiannual Yield to maturity (r) = 4%/2 = 2%

Years to maturity = 15

Coupon periods to maturity(n) = 15*2 = 30

Price of Bond can be computed with following equation:

P=C* (P/A,r, n) + F* (P/F,r, n)

P = C*\frac{(1+r)^n-1}{r*(1+r)^n}+F*\frac{1}{(1+r)^n}

putting the values:

(1 +0.02)30 - 1 P = 60 * 0.02 *(1 +0.02)302, (1 + 0.02)30

1.81136158410335 - 1 P = 60 * - 0.02 * 1.81136158410335 2,000 1.81136158410335

0.81136158410335 P = 60 * - 0.0362272316820671 +1104.14177795982

P = 60 * 22.3964555510044 + 1104.14177795982

P = 2447.92911102008

P= $2.447.93

Hope this will help, please do comment if you need any further explanation. Your feedback would be highly appreciated.